OpenStudy (drummmlers):

Put the following in slope intercept form: 1/3x + 1/4y = 1/2

OpenStudy (mathmale):

First: some necessary housecleaning. By 1/3x, do you mean (1/3)x or do you mean 1 / (3x)? too much ambiguity here. Retype your question with appropriate use of parentheses, and then I'd be happy to help you solve this problem.

OpenStudy (anonymous):

1/3x+1/4y=1/2 1/3x-1/2=-1/4y switch sides by addition and subtraction -1/3x+1/2=1/4 because we have a negative y we need to make it positive -4/3x+2/3=y divide 1/4 out of y and do it to both sides

OpenStudy (drummmlers):

\[\frac{ 1}{ 3 }x + \frac{ 1 }{ 4 }y = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

anything else you need help with?

OpenStudy (solomonzelman):

everything times 4, subtract 4/3x from both sides

OpenStudy (texaschic101):

slope intercept is y = mx + b 1/3x + 1/4y = 1/2 ---- multiply everything by the LCD (12) to get rid of fractions 4x + 3y = 6 -- subtract 4x from both sides 3y = -4x + 6 -- divide by 3 y = -4/3x + 2

OpenStudy (mathmale):

@drummmlers: Thank you. SO much clearer now!! I'd suggest that you identify and use the LCD first of all, to eliminate those fractions. This is of course not the only way to do this problem. Note that this is the approach used by @texaschic101.

OpenStudy (texaschic101):

I do not like dealing with fractions so that is why I do that. mathmale is correct in saying there is other ways to do this....